OpenStudy (anonymous):

? Which of the following is the simplified form of x plus 7 over 2 x plus 12 plus the fraction 6 over x squared minus 36 ? x plus 5 over 2 times the product x minus 6 x minus 5 over 2 times the product x plus 6 x minus 5 over 2 times the product x minus 6 x plus 5 over 2 times the product x plus 6

OpenStudy (anonymous):

it makes it easier to read if you type out the equations how they look instead of writing out the words.. ahah like:\[\frac{ x+7 }{ 2x +12}+\frac{ 6 }{ x^{2}-36 }\] the way you write out the words makes it confusing since they could actually mean totally different things, i'm just assuming this is what you meant

OpenStudy (anonymous):

so you want to factor the denominators first, then find a least common denominator so you can add the numerators together. then just simplify

OpenStudy (anonymous):

Okay and yes that's what meant

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

What do you need? :P

OpenStudy (luigi0210):

Did you get common denominators?

OpenStudy (anonymous):

No i'm confused... my friend tried explaining it differently now i'm all math out confused. D;

OpenStudy (luigi0210):

Can factor the bottoms?

OpenStudy (luigi0210):

\[\frac{ x+7 }{ 2(x+6) }+\frac{ 6 }{ (x-6)(x+6) }\]

OpenStudy (anonymous):

okay I see, but do we cross out the bottom part or?

OpenStudy (luigi0210):

Nope, nothing cancels. get common denominators

OpenStudy (anonymous):

so how do I do that, would it just be 6?

OpenStudy (luigi0210):

:l

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

I'm sorry, D:

OpenStudy (anonymous):

is it this \[\frac{ x+7 }{ 2x +12}+\frac{ 6 }{ x^{2}-36 }\]?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

oh i see @Luigi0210 got it, nvm

OpenStudy (anonymous):

?

OpenStudy (luigi0210):

Sorry about that.. but yea, get common denominators.. make the bottom the same for both expressions.

OpenStudy (anonymous):

so its not 6

OpenStudy (anonymous):

factor as shown to get \[\frac{ x+7 }{ 2(x+6) }+\frac{ 6 }{ (x-6)(x+6) }\] then the least common denominator will be \[2(x+6)(x-6)\]

OpenStudy (luigi0210):

Unless my math skills are gone

OpenStudy (anonymous):

so that will be 2x+12 x's 2x-12?

OpenStudy (anonymous):

then you combine like terms?

OpenStudy (anonymous):

\[\frac{ x+7 }{ 2(x+6) }+\frac{ 6 }{ (x-6)(x+6) }\] \[\frac{ x+7 }{ 2(x+6) }\times \frac{x-6}{x-6}+\frac{ 6 }{ (x-6)(x+6) }\times \frac{2}{2}\] \[\frac{(x+7)(x-6)}{2(x+6)(x-6)}+\frac{12}{2(x-6)(x+6)}\] \[=\frac{(x+7)(x-6)+12}{2(x-6)(x+6)}\]

OpenStudy (anonymous):

I'm lost, I'm sorry.

OpenStudy (anonymous):

yeah this is a bit annoying, but it is how you add fractions

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

suppose we replace \(x\) by 11

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

so \(x+6=11+6=17\) and \(x-6=11-6=5\) and also \((x+6)(x-6)=11\times 5=55\) and \(x+7=11+7=18\) and finally \(2(x+6)=2(11+6)=34\) then how would you add

OpenStudy (anonymous):

\[\frac{18}{34}+\frac{6}{55}\]

OpenStudy (anonymous):

you have to find a least common denominator for these fractions

OpenStudy (anonymous):

Okay,

OpenStudy (anonymous):

DAMN

OpenStudy (anonymous):

i made a mistake \(11+6=17\) and \(11-6=5\) so \((x+6)(x-6)=5\times 17=85\) and you have to add \[\frac{18}{34}+\frac{6}{85}\] that is what i meant

OpenStudy (anonymous):

how do you do it? you have to find a least common multiple of \(34\) and \(85\)

OpenStudy (luigi0210):

uhm, whats happening satellite?

OpenStudy (anonymous):

and to do that, you need to factor these numbers \(34=2\times 17\) and \(85=5\times 17\) so the least common denominator will be \[2\times 17\times 5\]

OpenStudy (anonymous):

then to add you would do this \[\frac{18}{34}\times \frac{5}{5}\] and \[\frac{6}{85}\times \frac{2}{2}\] that would make the denominators the same, so you can add

OpenStudy (anonymous):

\[\frac{18\times 5}{170}+\frac{6\times 2}{170}=\frac{18\times 5+6\times 2}{170}\]

OpenStudy (anonymous):

it is exactly the same with variables. you find a least common multiple of the denominators, and then multiply each fraction top and bottom by what you need to put both fractions with the same denominator

OpenStudy (anonymous):

for \(34\) and \(85\) you needed a 2, a 5 and a 17 for\(2(x+6)\) and \((x+6)(x-6)\) you need a \(2\) , a \(x+6\) and a \(x-6\)

OpenStudy (anonymous):

that is my explanation of this line \[\frac{ x+7 }{ 2(x+6) }\times \frac{x-6}{x-6}+\frac{ 6 }{ (x-6)(x+6) }\times \frac{2}{2}\] you are putting each fraction over the same denominator, so you can add don't let the stupid \(x\) cloud the idea, it is just the same as with numbers

OpenStudy (anonymous):

Sorrrrrrrrrrry, I left and forgot to say ill brb... D:

OpenStudy (anonymous):

@satellite73